OpenStudy (anonymous):

The sum of two numbers is 35 and their difference is 7. Find the numbers.

OpenStudy (nathanjhw):

21 and 14.

OpenStudy (thatonegirl_):

x+y=35 x-y=7 2x=42 x=21. 21+y=35 y=14